OpenStudy (anonymous):

four women must cross a bridge over a deep ravine in enemy territory in the middle of the night. The bridge will hold only two women at a time & it is necessary to carry a lantern while crossing. Persons must stay together while crossing. One of the women requires 5 minutes to cross, one takes 10 minutes, the third takes 20 minutes and the last takes 25 minutes (each can walk slower if necessary but no faster). Unfortunately, they have only one lantern amongh them. How can they make the crossing if they have only 60 minutes before the bridge is destroyed.

OpenStudy (anonymous):

4 women must cross a bridge over a deep ravine in enemy territory in the middle of the night. The bridge will only old two women at once and it is necessary to carry a lantern while crossing. One of the women requires 25 minutes, one 20 minutes, one 10 minutes and one 5 minutes. Each of them can walk slower but not faster. Unfortunately they have only one lantern among them. How can they make the crossing if they have only 60 minutes before the bridge is destroyed?

OpenStudy (anonymous):

first the 5 and 10 minute people cross, then the 5 minute person goes back with latern ( so 15minutes is used ) and you have 5,20 and 25 still to cross then the 20 and 25 minute goes across ( which means 25 +15 = 40mins is used ) then 10mintue person ( which has already crossed ) , goes back to the start with the latern ( which means 40+10 =50mintues used ). then the 5 and 10mintue cross ( 10+50 =60mintues ) and all have crossed .

OpenStudy (anonymous):

^ in one post

OpenStudy (anonymous):

Thank you!!!